# Detection of evolutionary shifts in variance under an Ornstein–Uhlenbeck model

### v1

@@ 4 Simulations
We conduct simulations to test the effectiveness of our method to detect the true shifts in the variance without losing too much accuracy in detecting the shifts in the optimal value. Results of Bayesian approaches are deeply influenced by prior distributions and the computation cost is relatively high. Therefore, we only compare our results with frequentist approaches in this paper. We compare our method with l1ou+pBIC, l1ou+BIC, and phyloEM in the simulations (Khabbazian, Kriebel, Rohe, and Ané, 2016; Bastide, Mariadassou, and Robin, 2017). For the measurement of performance, we consider the True positive and False positive rates for detecting the shifts in variance and optimal values. However, in shift detection, model identification problems can occur. A surrogate shift is a shift on the branch near the true shift (for example, shift on the branch of the parent node or the child node). It influences similar species as the true shift. Selecting a surrogate shift of the true shift could give better inference and prediction than selecting no shift. However, simply counting true positives and false positives might not give sufficient credit for selecting the surrogate shift. Therefore, we also use predictive log-likelihood difference from the true model to evaluate the method performances. For each simulation, we generate $1000$ test datasets and $100$ training datasets. We calculate the mean of log likelihood values over $1000$ test datasets using the estimated shifts from training sets. When the results of a method give higher predictive log-likelihood value, it indicates that the method performs better at predicting the trait values. In order to compare the results across different scenarios, we also calcualte the predictive log-likelihood of the true model, and finally calculate the difference in predictive log-likelihood between each method and the true model. Smaller predictive log-likelihood difference shows smaller gap between the results and the true model, indicating better prediction performance. We use OU model with fixed root to calculate the likelihood.
We simulate datasets under OU models along the 100-taxon Anolis lizards tree (Mahler, Ingram, Revell, and Losos, 2013). We consider 3 different scenario settings; only shifts in variance, only shifts in optimal value and both shifts in optimal value and shifts in variance. For each parameter setting, 100 simulations are conducted. For all simulations, we set $\alpha=1$ and $\sigma_{0}^{2}=2$. Since our method does not support the estimation of $\alpha$, we assume that the correct $\alpha$ value is known for all the methods.
4.1 Case I: Only shifts in variance
The first scenario setting is that only shifts in variance exist in the OU model — the optimal value is constant throughout the tree. We perform a number of simulation studies with a single shift in variance, and compare the detection results when the shift occurs in different positions.
The left figure of Figure 3 shows the different shift positions we set. The right figure of Figure 3 shows the true positive and false positive rates for each shift position. We see that shifts in positions near the root are easier to detect, which is consistent with our analysis in Section 2. When the shift contains a sufficient number of tips and the signal size is large enough, our method has the ability to detect the true shift.
Figure 3: The true positive v.s. false positive rates of our method when there are only shifts in variance
Since there is no true shift in optimal value in this scenario, we here compare the false positive shifts detected in optimal values from different methods. From Figure 4, when the shift in variance has a sufficient number of tips below it (Shift 195) and the shift size is large enough, other methods will detect many false positives in optimal values by misinterpreting the signal as shifts in optimal value. However, our method can successfully detect the true shifts in variance in these cases, so that the false positive numbers for shifts in optimal value can be effectively controlled. The plots of other shift settings show similar patterns and are in the supplemental materials.
Figure 4: The false positive rates of different methods for detecting shifts in optimal value when there are only shifts in variance
Finally, We compare the predictive log-likelihood difference with the true model for each method. Figure 5 shows that the predictive log likelihood difference is smaller for our method than other methods, especially when the signal size is large and the shifts are near the root. This shows that our method has better predictive ability than other methods when there are only shifts in variance.
Figure 5: The predictive log-likelihood difference of different methods when there are only shifts in variance
4.2 Case II: Only shifts in optimal value
Simulations are conducted under scenarios with 1, 3, 7, or 12 shifts. The results are shown in Figures 67 and 8. From Figure 6, we see that ShiVa shows no significant weakness in detecting the true shifts in variance. Furthermore, ShiVa generally detects fewer false positives even when there are no shifts in optimal value to mislead the other methods. We also check the number of false positive shifts in variance under our method. Figure 7 shows that the number of false positives is limited. By comparing the predictive log-likelihood difference with true models, we find that our method performs worse than other methods when the number of shifts is large and the shift sizes are large. However, the gap between our method and the other methods is not substantial.
Figure 6: The true positive v.s. false positive rates of different methods when there are only shifts in optimal value
Figure 7: The false positive rates of the new method for shifts in variance when there are only shifts in optimal value
Figure 8: The predictive log-likelihood difference of different methods when there are only shifts in optimal value
4.3 Case III: Both shifts in optimal value and shifts in variance
We simulate one shift in optimal value and one shift in variance. We consider three different relationships between the shift in optimal value and the shift in variance: the shifts are on independant lineages; the shift in variance is ancestral to the shift in optimal value; and the shift in optimal value is ancestral to the shift in variance. The simulated shift configurations are shown in Figure 9.
Figure 10 shows the true positive and false positive rate for each method in each scenario for detecting shifts in optimal value. Overall, our method has much fewer false positives for shifts in optimal value than other methods. All the other methods have very high false positive rates since they wrongly detect the shift in variance as many different shifts in the optimal value. In addition to controlling the false positive rates, ShiVa performs generally well in terms of detecting the true positive shifts in optimal value except for limited cases. Figure 12 compares the predictive log-likelihood for each method in these scenarios. We see that our method has the smallest predictive log-likelihood difference from the true model. This shows that when there are shifts in variance, even if there are also shifts in optimal value, ShiVa selects better predictive models than methods that only select shifts in optimal value.
Figure 11 shows the true positive and false positive rate of ShiVa for shifts in variance. In order to show how the magnitudes of shift in variance and shift in optimal value influence the shift detection results, we keep the magnitude of shift in optimal value as 10 and vary magnitude of shift in variance in the left plot; and we keep the magnitude of shift in variance as 20 and vary magnitude of shift in optimal value in the right plot. It shows that when the shift in optimal value is under the shift in variance, the shift in variance is easiest to detect. When the shift in variance is under the shift in mean, the shift in variance is hardest to detect. Furthermore, when the signal size of shift in variance is much larger than the signal size of shift in optimal value, ShiVa has better performance with higher true positive rates and lower false positive rates.
Figure 9: Shift configurations of the simulations. Circle: shift in optimal value; Rectangle: shift in variance
Figure 10: The true positive v.s. false positive rates of different methods for shifts in mean when there are both shifts in variance and optimal value
Figure 11: The true positive v.s. false positive rates of ShiVa for shifts in variance when there are both shifts in variance and optimal value. Left: Vary magnitude of shift in variance. Right: Vary magnitude of shift in optimal value
Figure 12: The predictive log-likelihood difference of different methods when there are both shifts in variance and optimal value
4.4 Discussion about assumption violations
@@
In Section 2, we noted that our assumptions about the shifts in variance—at most one shift will occur on any branch; and the shifts occur at the beginning of the branches—are restrictions on the space of possible models, so in particular if these assumptions are not satisfied, then the model is misspecified. In this subsection, we conduct simulations in which these assumptions are violated.
@@
Firstly, we conduct a series of simulations with the shift occuring in different positions along a fixed branch. The different locations of shift are shown in Figure 13 (left). Figure 14 shows the false positive v.s. true positive rates and the predictive log-likelihood difference of the new method with shifts in different locations. Location here is a parameter ranging from 0 to 1, with 0 meaning the beginning of the branch and 1 meaning the end of the branch. The ability of the method to detect the shift in variance is not greatly impacted by the position of the shift along the branch. Furthermore, when the shift size is small, the log-likelihood of the model is not much reduced in situations where the location of the shift is not at the beginning of the branch. However, when the magnitude of the shift is larger, the model misspecification does cause a substantial decrease in log-likelihood when the shift is not located at the begining of the branch. Overall, our method is robust to violations of our assumption that shifts occur at the beginning of a branch.
@@ 5 Case study: Cordylid data
Cordylidae is a family of lizards in Sub-Sarharan Africa (Stanley, Bauer, Jackman, Branch, and Mouton, 2011). Broeckhoven et al. (2016) selected 28 species of Cordylidae, including five outgroup taxa, and reconstructed the phylogenetic tree based on their DNA sequences using BEAST. They also collected 20 morphological traits, including spine, limb and head dimensions and used pPCA (Revell, 2009) to perform dimension reduction. Based on their analysis, the first principal component pPCA1 represented a gradient from highly armored species to lightly armored species. The phylogenetic tree and principal component data are available in the R package phytools: http://www.phytools.org/Rbook/. We used the cordylid tree and the first principal component they provided to compare the shift detection results of $\ell$1ou, phyloEM and ShiVa.
Figure 15 shows the results given by different methods. ShiVa uses cross validation to tune the hyperparameter $\lambda_{1}$, therefore the final results have randomness. We suggest to repeat it 3 times and select the best model based on BIC in real data analysis. It shows that the model estimated by ShiVa gives the highest log-likelihood (18.2) and the lowest BIC (40.24). $\ell$1ou+pBIC and phyloEM detected limited numbers of shifts, with lower log-likelihood and higher BIC compared to the other two methods. Comparing the results of $\ell$1ou+BIC and ShiVa, $\ell$1ou+BIC detected branches 28, 43, 44 as shifts in optimal value, and ShiVa detected a shift in variance: branch 47, without detecting the above shifts in optimal value. From the results of ShiVa, it is possible that the subgroup Cordylus has a different variance from other taxa. And this interesting finding may be neglected, and indeed detected as multiple false shifts in optimal values using the other methods.
@@
Figure 15: The shift detection results of cordylid data of $\ell$1ou+pBIC (left top), $\ell$1ou+BIC (right top), phyloEM (left bottom), ShiVa (right bottom)
@@ 6 Conclusion
We performed simulations of three different situations. When there are shifts in variance, our method has much lower false positive rate for shifts in mean than methods which do not model shifts in variance, since other methods wrongly detect the shift in variance as a large number of shifts in mean. As a result, our method has higher predictive log-likelihood than methods which do not model shifts in variance. When there are only shifts in optimal value, our method performs slightly worse than other methods when the number of shifts is large and the shift sizes are large. However, the gap between our method and the other methods is not substantial. Therefore, when shifts in variance exist, our method can effectively detect the true shifts or surrogates of the shifts in variance, leading to more accurate detection of shifts in optimal value and better predictive ability. And when there are no shifts in variance, our method has similar performance to other methods.
@@
We applied our method to the first principal component of trait values for the cordylid data from Broeckhoven et al. (2016). ShiVa detected 7 shifts in optimal value and 3 shifts in variance and the model given by ShiVa has the highest log-likelihood and lowest BIC compared to $\ell$1ou and phyloEM. The validation result also shows Shiva would be unlikely to estimate the model it did if any of the models selected by other methods were true.
@@
For model simplicity, we assume that the variance parameter is constant along each branch and that all shifts in variance happen at the beginning of the branch. Simulations of cases where these assumptions are violated show that our method is robust to misspecification in these assumptions. An interesting future research direction is to extend our model to allow shifts in variance at internal postions along a branch. This would allow us to estimate the exact time of a shift in variance. Writing the likelihood for this case is straightforward, but more work may be needed to ensure stability of the estimates. Another direction for future work is to improve the computational efficiency of the method. The likelihood calculation involves a large number of matrix inverse computations,which can be computationally expensive for large trees. Bastide et al. (2021) provide an efficient algorithm to calculate the log likelihood and its derivatives for certain phylogenetic models. If this method could be adapted to our method, it would lead to a major improvement in the computational speed, allowing our method to scale to larger phylogenies.
7 Data accessibility

### v2

@@ 4 Simulations
4.1 Comparison of performance
In the main simulation settings, we set $\alpha=1$ and the original variance $\sigma^{2}_{0}=2$. These parameter choices represent a moderate level of phylogenetic signal and baseline trait variance. For each parameter setting, we conducted 50 simulations. To assess the robustness of our conclusions, we also conducted additional simulations with different values of $\alpha$, which are provided in the supplementary materials. The results of these additional simulations are consistent with the main findings presented here. In the following simulations, $\alpha$ is assumed to be unknown. For ShiVa, we obtain $\alpha$ by fitting a null model (without shifts) using the phylolm package.
We compared the performance of different methods based on three key metrics: True Positives (TP), False Positives (FP), predictive log-likelihood, and computational time. Here, True Positives (TP) refer to the number of true shifts correctly detected, while False Positives (FP) refer to the number of false shifts incorrectly detected. However, the positions of shifts are not always identifiable—for example, under a 3-branch tree, any two shifts can lead to an identical model. In such cases, TP and FP may not be reliable metrics, as they would count the true model as including one false positive and missing one true positive, even though the models are effectively equivalent. Figure 3 illustrates the TP and FP results across the various methods under different scenarios. When there is only a shift in the optimal value, ShiVa achieves a well-balanced trade-off between True Positives and False Positives. It effectively detects shifts in the optimal value without generating excessive False Positives. The performance of ShiVa is comparable to other methods that only consider shifts in the optimal value even if only shift in optimal value is present. In contrast, PCMFit produces a much higher number of False Positives, particularly when the signal size is small.
In the case of shifts in variance, ShiVa performs well in detecting variance shifts effectively while maintaining control over False Positives. Although PCMFit shows a higher True Positive rate compared to ShiVa, it comes at the cost of generating significantly more False Positives. Methods like $\ell$1ou and phyloEM, which are designed to detect shifts in optimal value, face challenges when a shift in variance is present. As the magnitude of the variance shift increases, these methods incorrectly interpret the variance signal as multiple shifts in the optimal value. Consequently, their False Positives increase substantially. In contrast, ShiVa successfully distinguishes between variance and mean shifts, effectively controlling the False Positive rate for optimal value shifts, even as the variance shift signal strengthens.
When both shifts in optimal value and variance are present, ShiVa continues to perform competitively. Although it detects slightly fewer True Positives compared to PCMFit, its False Positive rate is much lower, indicating a better balance between sensitivity and specificity. PCMFit, while capable of identifying more shifts overall, tends to produce excessive False Positives. This leads to reduced reliability in shift detection, especially in noisy data.
(a) 1 shift in optimal value
(b) 1 shift in variance
(c) 1 shift in optimal value + 1 shift in variance (fixed gamma)
(d) 1 shift in optimal value + 1 shift in variance (fixed beta)
Figure 3: Comparison of True positives and False positives of different methods
In addition to TP and FP, we also evaluated predictive log-likelihood to assess the prediction accuracy of the models estimated by different methods. Regarding identifiability, the OU model is identifiable when no shifts are present. However, previous studies have shown that the exact number and locations of shifts along an edge cannot always be fully determined due to identifiability issues. For example, any two of the three branches connected to a node can yield equivalent models. Even if the selected shifts do not match the true model exactly, choosing a close surrogate shift may be preferable to missing the shift entirely. In such cases, the true positive versus false positive framework might misrepresent performance: a method selecting a surrogate shift is penalized with a false positive, even though it provides a more reasonable approximation than failing to detect the shift. Predictive log-likelihood addresses this issue by evaluating how well the selected models predict new data. For each simulation, we generate 1,000 test datasets for each of the 2,000 simulated datasets, using the same tree and shift configurations. We then compute the mean log-likelihood across the 1,000 test datasets based on the estimated shifts from the training sets. Figure 4 illustrates the differences in predictive log-likelihood between the estimated models and the true model. When shifts occur only in optimal value, ShiVa performs comparably to $\ell$1ou and phyloEM, demonstrating similar predictive precision. However, when shifts in variance are introduced, ShiVa shows a notable improvement, achieving higher predictive log-likelihood than the other methods. In contrast, PCMFit consistently exhibits the lowest predictive log-likelihood across all scenarios, likely due to overfitting.
(a) 1 shift in optimal value
(b) 1 shift in variance
(c) 1 shift in optimal value + 1 shift in variance (fixed gamma)
(d) 1 shift in optimal value + 1 shift in variance (fixed beta)
Figure 4: Comparison of predictive log-likelihood of different methods
Finally, we compare the average computational time across different methods. Methods that only account for shifts in optimal values, such as $\ell$1ou, are the fastest, with $\ell$1ou being particularly efficient due to its use of the LASSO path method. ShiVa performs similarly to phyloEM but is slightly slower in comparison. On the other hand, PCMFit requires significantly more time due to its handling of complex computations, especially when both shifts in optimal values and variance are present. The current PCMFit implementation allows parallel computing to partially compensate for the disadvantage of the method in terms of computational complexity. For the comparisons in Table 1, we ran all methods single-threaded to provide a fairer comparison of computation required by all methods.
Table 1: Computational Time Comparison for Different Methods
Average Computation Time Shift Position (mean) Beta Shift Position (variance) Gamma ShiVa l1ou-pBIC l1ou-BIC phyloEM PCMFit 71 1 0 0 95.38 8.09 8.31 68.15 11484.33 71 3 0 0 77.39 8.29 8.57 70.10 13592.56 71 5 0 0 73.86 8.59 8.78 69.92 12227.06 71 -1 0 0 143.39 7.86 8.24 67.81 12890.20 71 -3 0 0 75.70 8.31 8.45 70.02 16215.21 71 -5 0 0 70.33 8.37 8.57 69.35 11962.69 0 0 195 1 70.22 8.08 8.30 68.63 15171.28 0 0 195 1.5 83.31 8.08 8.42 67.72 15680.57 0 0 195 2 77.95 8.04 8.32 70.84 20622.02 0 0 195 3 87.73 8.52 8.81 70.93 16349.04 0 0 195 5 90.05 8.10 8.34 67.70 17358.66 0 0 195 7 97.68 8.18 8.53 67.85 19644.00 0 0 195 -1 101.92 8.04 8.30 67.95 15269.57 0 0 195 -1.5 102.83 7.58 7.74 65.42 15557.74 71 5 195 5 79.32 8.30 8.68 68.18 22836.10 71 3 195 5 129.45 8.55 8.73 68.68 24499.03 71 1 195 5 75.43 8.26 8.92 69.12 18023.39 71 -1 195 5 85.47 7.93 8.24 69.71 19622.12 71 -3 195 5 68.37 7.88 8.51 69.10 24580.25 71 -5 195 5 77.10 7.89 8.21 70.45 23177.43 71 5 195 1 76.73 8.23 8.53 70.06 16574.76 71 5 195 1.5 72.93 8.27 8.59 69.33 13153.55 71 5 195 2 76.58 8.18 8.61 68.71 15742.43 71 5 195 3 90.45 8.17 8.54 69.52 22800.45 71 5 195 7 87.67 8.42 8.67 69.44 23418.19 71 5 195 -1 92.83 8.23 8.45 70.31 16642.62 71 5 195 -1.5 95.15 8.31 8.51 70.79 17517.06
4.2 Relationship between shift position and detection difficulty
To investigate the difficulty of detecting shifts at specific positions on a phylogenetic tree, we conducted simulations for every internal branch of the tree. For each branch, we performed 50 simulations where a shift in the optimal value occurred and 50 simulations where a shift in variance occurred. We then calculated the detection probabilities for each shift using ShiVa, considering different positions on the tree. Figure 5 presents the detection probabilities for various branches, highlighting the different detection patterns for shifts in optima versus shifts in variance.
The results indicate that shifts in optima and shifts in variance exhibit distinct detection patterns. For shifts in the optimal value, those occurring near the root or the tips are easier to detect compared to shifts occurring in the middle regions of the tree. In contrast, shifts in variance are generally harder to detect. Our findings show that shifts in variance are more likely to be detected when they occur on branches with a larger number of descendant tips and longer branch lengths, which provides a stronger signal for the variance change. This pattern aligns with our analysis in the Evolutionary shifts in optimal value and variance subsection. The overall results emphasize the importance of the tree structure and branch characteristics in determining the detectability of evolutionary shifts, with shifts in variance requiring more stringent conditions for reliable detection compared to shifts in optima.
Figure 5: The detection probability of a shift on different branches (ShiVa)
4.3 The influence of estimation of $\alpha$
We conducted simulations to investigate how the misestimation of $\alpha$ affects ShiVa’s shift detection results. In these simulations, we set the true value of $\alpha$ to 1 (corresponding to $\log(\alpha)=0$). To assess the impact of different levels of misestimation, we tested a range of estimated $\alpha$ values, using $\log(\alpha)=-5,-3,-1,0,1,3,5$ as inputs to the ShiVa function. The detection results were evaluated by comparing the predictive log-likelihoods across these different estimated values of $\alpha$.
For this analysis, we focused on scenarios with one shift in the optimal value and one shift in variance, representing a typical use case for ShiVa. The results suggest that, in most cases, moderate misestimation of $\alpha$ has little effect on ShiVa’s performance. However, the impact becomes more pronounced in extreme cases — particularly when the estimated $\alpha$ is very large (e.g., $\exp(5)$) and the shift signal size is also very large. This indicates that while ShiVa is generally robust to $\alpha$ misestimation, severe overestimation combined with strong signals may reduce detection accuracy. Overall, this analysis supports the practical simplification of using a reasonable ad-hoc estimate of $\alpha$ in applications.
(a) 1 shift in optimal value: $\beta$ = 3
(b) 1 shift in variance: $\gamma$ = 3
Figure 6: Comparison of predictive log-likelihood of ShiVa with different estimation of $\alpha$
4.4 The effect of shift position within a branch
@@
In Trait evolution with shifts in both optimal value and variance, we noted that our assumptions about the shifts in variance—at most one shift will occur on any branch; and the shifts occur at the beginning of the branches—are restrictions on the space of possible models, so in particular if these assumptions are not satisfied, then the model is misspecified. In this subsection, we conduct simulations in which these assumptions are violated.
@@
Firstly, we conduct a series of simulations with the shift occuring in different positions along a fixed branch. The different locations of the shift are shown in Figure 7 (left). Figure 8 shows the True Positives v.s. False Positives rates and the predictive log-likelihood difference between ShiVa and the true model with shifts in different locations. Location here is a parameter ranging from 0 to 1, with 0 meaning the beginning of the branch and 1 meaning the end of the branch. The ability of the method to detect the shift in variance is not greatly impacted by the position of the shift along the branch. Furthermore, when the shift size is small, the log-likelihood of the model is not much reduced in situations where the location of the shift is not at the beginning of the branch. However, when the magnitude of the shift is larger, the model misspecification does cause a substantial decrease in log-likelihood when the shift is not located at the begining of the branch. Overall, our method is robust to violations of our assumption that shifts occur at the beginning of a branch.
@@ 5 Case study: Cordylid data
Cordylidae is a family of lizards in Sub-Sarharan Africa (Stanley, Bauer, Jackman, Branch, and Mouton, 2011). Broeckhoven et al. (2016a) selected 28 species of Cordylidae, including five outgroup taxa, and reconstructed the phylogenetic tree based on their DNA sequences using BEAST. They also collected 20 morphological traits, including spine, limb and head dimensions and used pPCA (Revell, 2009) to perform dimension reduction. Based on their analysis, the first principal component pPCA1 represented a gradient from highly armored species to lightly armored species. The phylogenetic tree and principal component data are available in the R package phytools: http://www.phytools.org/Rbook/. We used the cordylid tree and the first principal component they provided to compare the shift detection results of $\ell$1ou, phyloEM and ShiVa.
Figure 9 shows the results given by different methods. ShiVa uses cross validation to tune the hyperparameter $\lambda_{1}$, therefore the final results have randomness. We repeat the method 3 times and select the best model based on BIC for this real data analysis. The model estimated by ShiVa gives the highest log-likelihood (18.2) and the lowest BIC (40.24). $\ell$1ou+pBIC and phyloEM detected limited numbers of shifts, with lower log-likelihood and higher BIC compared to the other two methods. Comparing the results of $\ell$1ou+BIC and ShiVa, $\ell$1ou+BIC detected branches 28, 43, 44 as shifts in optimal value, and ShiVa detected a shift in variance: branch 47, without detecting the above shifts in optimal value. From the results of ShiVa, it is possible that the subgroup Cordylus has a different variance from other taxa. This interesting finding may be neglected, and indeed detected as multiple false shifts in optimal values using the other methods.
@@
Figure 9: The shift detection results of cordylid data of $\ell$1ou+pBIC (left top), $\ell$1ou+BIC (right top), phyloEM (left bottom) and ShiVa (right bottom)
@@ 6 Conclusion
Our results showed that ShiVa effectively balances True Positives (TP) and False Positives (FP) across different scenarios. When there was only a shift in the optimal value, ShiVa performed comparably to other methods, such as $\ell$1ou and phyloEM, maintaining similar predictive accuracy while achieving a better balance between TP and FP compared to PCMFit, which tended to overfit in weak signal scenarios. When a shift in variance was present, ShiVa demonstrated a significant advantage over other methods, achieving higher predictive log-likelihood and effectively distinguishing between variance and mean shifts. This led to fewer False Positives compared to $\ell$1ou and phyloEM, which often misinterpreted variance shifts as multiple shifts in optimal value. Although PCMFit achieved higher TP rates, it generated significantly more FP, indicating overfitting.
@@
We applied our method to the first principal component of trait values for the cordylid data from Broeckhoven et al. (2016a). ShiVa detected 7 shifts in optimal value and 3 shifts in variance and the model given by ShiVa has the highest log-likelihood and lowest BIC compared to $\ell$1ou and phyloEM. The validation result also shows that Shiva would be unlikely to estimate the model it did if any of the models selected by other methods were true.
@@
For model simplicity, we assume that the variance parameter is constant along each branch and that all shifts in variance happen at the beginning of the branch. Simulations of cases where these assumptions are violated show that our method is robust to misspecification in these assumptions. An interesting future research direction is to extend our model to allow shifts in variance at internal positions along a branch. This would allow us to estimate the exact time of a shift in variance. Writing the likelihood for this case is straightforward, but more work may be needed to ensure stability of the estimates. Another direction for future work is to improve the computational efficiency of the method. The likelihood calculation involves a large number of matrix inverse computations,which can be computationally expensive for large trees. Bastide et al. (2021) provide an efficient algorithm to calculate the log likelihood and its derivatives for certain phylogenetic models. If this method could be adapted to our method, it would lead to a major improvement in the computational speed, allowing our method to scale to larger phylogenies.
Appendix A Supplemental Materials
f
(a) 3 shifts in optimal value
(b) 7 shifts in optimal value
(c) 1 shift in variance (different position)
(d) 2 shifts in variance
Figure 12: Comparison of True positives and False positives of different methods (Supplementary simulations)
(a) 3 shifts in optimal value
(b) 7 shifts in optimal value
(c) 1 shift in variance (different position)
(d) 2 shifts in variance
Figure 13: Comparison of predictive log-likelihood (difference with true model) of different methods (Supplementary simulations)
Table 4: Computational Time Comparison for Different Methods (Supplementary simulations)
Average Computation Time Shift Position (mean) Beta Shift Position (variance) Gamma ShiVa l1ou-pBIC l1ou-BIC phyloEM PCMFit 43,71,192 1 0 0 86.80 7.66 8.43 69.64 11561.58 43,71,192 3 0 0 74.95 8.37 8.36 68.09 27477.04 43,71,192 5 0 0 88.87 8.76 8.86 68.05 36900.89 43,71,192 -1 0 0 71.88 7.99 8.40 69.21 13952.25 43,71,192 -3 0 0 67.89 8.61 8.56 70.58 35867.14 43,71,192 -5 0 0 76.54 9.13 9.16 72.69 40202.81 6,71,197,3,88,191,98 1 0 0 117.94 7.80 8.05 70.53 14863.06 6,71,197,3,88,191,98 3 0 0 80.94 6.06 6.08 70.63 31135.83 6,71,197,3,88,191,98 5 0 0 81.68 8.55 8.86 67.94 59322.66 6,71,197,3,88,191,98 -1 0 0 83.20 7.79 8.04 69.02 15652.61 6,71,197,3,88,191,98 -3 0 0 98.53 7.42 6.84 69.47 29763.69 6,71,197,3,88,191,98 -5 0 0 69.55 7.99 7.85 70.15 60076.49 0 0 196 1 68.06 8.08 8.35 69.46 14515.10 0 0 196 1.5 72.20 8.03 8.31 69.51 15649.65 0 0 196 2 83.65 8.39 8.54 70.18 17266.36 0 0 196 3 90.27 7.61 8.14 69.24 18907.37 0 0 196 5 89.08 8.27 8.42 69.61 18364.71 0 0 196 7 101.44 8.03 8.51 67.09 18858.03 0 0 196 -1 99.09 8.26 8.66 67.89 16773.86 0 0 196 -1.5 113.54 8.21 8.72 67.30 19419.61 0 0 195,130 1,-0.5 80.75 8.37 8.68 70.94 16275.18 0 0 195,130 3,-1 93.39 8.11 8.39 70.21 18297.14 0 0 195,130 5,-1.5 159.13 12.06 12.58 116.55 36602.57 0 0 195,130 7,-1.5 114.15 8.60 8.86 69.08 22939.48 0 0 195,130 1 85.99 8.15 8.31 70.17 15881.60 0 0 195,130 3 91.66 8.10 8.38 70.03 17969.21 0 0 195,130 5 104.25 8.69 9.17 71.10 25327.97 0 0 195,130 7 95.08 8.63 8.94 69.17 23545.64
Appendix B Data accessibility